OpenStudy (anonymous):

Find the slope and the y-intercept 3-1/4y=10x

OpenStudy (anonymous):

-40, 12

OpenStudy (anonymous):

so the slope is -40, and the y-intercept is (0,12)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

how did you simplify the expression?

OpenStudy (anonymous):

1/4y = 3-10x y = 12 - 40x

OpenStudy (anonymous):

oh rats.. i thot y was in the denominator.. so sorry

OpenStudy (anonymous):

x= the first day a ticket goes on sale